# Lightweight and Interpretable Transformer via Mixed Graph Algorithm Unrolling for Traffic Forecast

### v2

@@ A.1 Lightweight Learning Models
Given the importance of parameter reduction for increasingly large deep learning models, there exist various studies towards this common goal across application domains. For large language models (LLMs), smaller transformer models have been achieved via low-rank assumption (Hu et al., 2022) or parameter quantization (Frantar et al., 2023). For speech enhancement, Oostermeijer et al. (2021) proposes local causal self-attention, while Zhao & Madhu (2025) proposes parameter sharing across spectral and time dimensions. For low-light image enhancement, Brateanu et al. (2025) processes the Y-channel via convolution, pooling, and multi-head self-attention, separately from the U- and V-channels. For hyperspectral image classification, Sun et al. (2024) proposes a novel multiscale 3D-2D mixed CNN design. For cervical cancer detection, Pacal (2024) substitutes larger MBConv blocks in the MaxViT architecture with smaller ConvNeXtv2 blocks, and MLP blocks with GRN-based MLPs. Similar architecture miniaturization techniques have been proposed for edge devices (Mishra & Gupta, 2024) and resource-constrained environments (Liu et al., 2024).
@@
Thuc et al. (2024) achieves a lightweight model for image interpolation in a more principled manner: first design a linear-time optimization algorithm minimizing a chosen undirected graph smoothness prior (such as graph Laplacian regularizer (GLR) (Pang & Cheung, 2017) or graph total variation (GTV) (Bai et al., 2019)), then unroll its iterations into neural layers to compose a “white-box” transformer-like neural net for data-driven parameter learning. This algorithm unrolling approach results in a lightweight and mathematically interpretable network. Our current work differs from Thuc et al. (2024) in that we model the more complex node-to-node relationships in spatial-temporal data using a mixed graph: i) an undirected graph ${\mathcal{G}}^{u}$ to capture spatial correlations among geographically near stations, and ii) a directed graph ${\mathcal{G}}^{d}$ to capture sequential relationships along the temporal dimension. In particular, the notion of “frequencies” for directed graphs is not well understood in the graph signal processing (GSP) community. We are the first to define this notion and promote low-pass signal reconstruction via our designed $\ell_{2}$-norm directed graph Laplacian regularizer (DGLR) variational term in eq. 2 and $\ell_{1}$-norm directed graph total variation (DGTV) variational term in eq. 3, both derived from the graph shift operator (GSO) (Chen et al., 2015). To the best of our knowledge, we are the first define and promote smooth signals on a directed graph in a data-driven transformer setting.
@@ A.2 Deep Models for Traffic Forecasting
Traffic forecasting problems contain spatial and temporal correlations between stations and time steps; thus, Graph Neural Network (GNN) is naturally employed in early research. Early research, such as STGCN (Yu et al., 2018), STSGCN (Song et al., 2020) and GGRNN (Ruiz et al., 2020), employs Graph Convolutional Networks (GCNs) to capture spatial relations and uses 1D convolutions and RNNs to process temporal relationships. ASTGNN (Ta et al., 2022), Graph WaveNet (Wu et al., 2019) and AGCRN (Bai et al., 2019) explored adaptive graph settings to capture potential non-local signal relations over the graph, trading in interpretability for improved performance. Graph Attention Networks (GATs) learn graph weights from embedded signals to enable flexible and data-targeted graph learning, typical examples include GMAN (Zheng et al., 2020) and ST-Wave (Fang et al., 2023). All GNNs above focus only on the spatial correlations, while the product graph (PG) is an intuitive model that describes local spatial-temporal relations at the same time. Einizade et al. (2024) presents a continuous product graph neural network for spatial-temporal forecasting, which achieves superior performance to the previous most-popular graph methods.
Transformers extend the powerful self-attention mechanism to all spatial-temporal data points, which enhances model performance with an extra-large number of parameters. Popular transformer-based models for traffic forecasting employ two transformers to process spatial and temporal dimensions separately (Xu et al., 2020; Chen et al., 2022; Huo et al., 2023), while state-of-the-art transformers introduce adaptive local attentions (Liu et al., 2023), time-delay (Jiang et al., 2023), or pre-training (Gao et al., 2024).
In recent years, simple linear models have been proposed to challenge the effectiveness and efficiency of heavy transformers. DLinear (Zeng et al., 2023) simply separates the input signals as trends and residuals, and employs 2 simple linear layers shared across nodes for prediction, but beats some of the transformer baselines in long-term forecasting. Other representative simple linear models include STID (Shao et al., 2022) and SimpleTM Chen et al. (2025).
@@
Appendix C Interpreting Linear System eq. 10
@@
${\mathbf{r}}^{\tau}\triangleq({\mathbf{L}}^{d}_{r})^{\top}(\frac{{\bm{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}.$ (54)
Solution ${\mathbf{x}}^{\tau+1}$ to linear system eq. 10 can now be written as
@@
Appendix D Deriving Solution to Optimization eq. 15
The optimization for $\boldsymbol{\phi}$} in objective eq. 15 can be performed entry-by-entry. Specifically, for the $i$-th entry $\phi_{i}$:
@@
Alternative solution with subgradients: Denote $g_{1}({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1},g_{2}({\boldsymbol{\phi}})=({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|_{2}^{2}$, the subgradient of $g({\boldsymbol{\phi}})$ is
$\partial g({\boldsymbol{\phi}})=\{{\mathbf{g}}+\nabla g_{2}({\boldsymbol{\phi}})|{\mathbf{g}}\in\partial g_{1}({\boldsymbol{\phi}})\}=\{\mu_{d,1}\sigma+{\bm{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})|\sigma\in\partial(\|{\boldsymbol{\phi}}\|_{1})\},$ (61)
@@
Appendix E Parameterizing Conjugated Gradient Method in Solving Linear Equations
@@
Algorithm 1 Conjugate Gradient (CG) Method to Solve ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$
@@
7:  $\boxed{\alpha_{k}\leftarrow\dfrac{{\mathbf{r}}_{k}^{\top}{\mathbf{r}}_{k}}{{\mathbf{p}}_{k}^{\top}{\mathbf{A}}{\mathbf{p}}_{k}}}$
@@
10:  $\boxed{\beta_{k}\leftarrow\dfrac{{\mathbf{r}}_{k+1}^{\top}{\mathbf{r}}_{k+1}}{{\mathbf{r}}_{k}^{\top}{\mathbf{r}}_{k}}}$

### v3

@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Lightweight Learning Models
Given the importance of parameter reduction for increasingly large deep learning models, there exist various studies towards this common goal across application domains. For large language models (LLMs), smaller transformer models have been achieved via low-rank assumption (Hu2022LoRA) or parameter quantization (Frantar2023GPTQ). For speech enhancement, oostermeijer2021lightweight proposes local causal self-attention, while zhao2025eusipco proposes parameter sharing across spectral and time dimensions. For low-light image enhancement, Brateanu2025 processes the Y-channel via convolution, pooling, and multi-head self-attention, separately from the U- and V-channels. For hyperspectral image classification, Sun2024 proposes a novel multiscale 3D-2D mixed CNN design. For cervical cancer detection, Pacal2024 substitutes larger MBConv blocks in the MaxViT architecture with smaller ConvNeXtv2 blocks, and MLP blocks with GRN-based MLPs. Similar architecture miniaturization techniques have been proposed for edge devices (Mishra2024) and resource-constrained environments (Liu2024).
@@
Do2024 achieves a lightweight model for image interpolation in a more principled manner: first design a linear-time optimization algorithm minimizing a chosen undirected graph smoothness prior (such as graph Laplacian regularizer (GLR) (pang17) or graph total variation (GTV) (bai19)), then unroll its iterations into neural layers to compose a “white-box” transformer-like neural net for data-driven parameter learning. This algorithm unrolling approach results in a lightweight and mathematically interpretable network. Our current work differs from Do2024 in that we model the more complex node-to-node relationships in spatial-temporal data using a mixed graph: i) an undirected graph ${\mathcal{G}}^{u}$ to capture spatial correlations among geographically near stations, and ii) a directed graph ${\mathcal{G}}^{d}$ to capture sequential relationships along the temporal dimension. In particular, the notion of “frequencies” for directed graphs is not well understood in the graph signal processing (GSP) community. We are the first to define this notion and promote low-pass signal reconstruction via our designed $\ell_{2}$-norm directed graph Laplacian regularizer (DGLR) variational term in Equation 2 and $\ell_{1}$-norm directed graph total variation (DGTV) variational term in Equation 3, both derived from the graph shift operator (GSO) (chen15). To the best of our knowledge, we are the first define and promote smooth signals on a directed graph in a data-driven transformer setting.
@@ A.2 Deep Models for Traffic Forecasting
Traffic forecasting problems contain spatial and temporal correlations between stations and time steps; thus, Graph Neural Network (GNN) is naturally employed in early research. Early research, such as STGCN (Yu_2018), STSGCN (song2020spatial) and GGRNN (ruiz2020gated), employs Graph Convolutional Networks (GCNs) to capture spatial relations and uses 1D convolutions and RNNs to process temporal relationships. ASTGNN (ta2022adaptive), Graph WaveNet (10.5555/3367243.3367303) and AGCRN (bai19) explored adaptive graph settings to capture potential non-local signal relations over the graph, trading in interpretability for improved performance. Graph Attention Networks (GATs) learn graph weights from embedded signals to enable flexible and data-targeted graph learning, typical examples include GMAN (zheng2020gman) and ST-Wave (fang2023spatio). All GNNs above focus only on the spatial correlations, while the product graph (PG) is an intuitive model that describes local spatial-temporal relations at the same time. einizade2024continuous presents a continuous product graph neural network for spatial-temporal forecasting, which achieves superior performance to the previous most-popular graph methods.
Transformers extend the powerful self-attention mechanism to all spatial-temporal data points, which enhances model performance with an extra-large number of parameters. Popular transformer-based models for traffic forecasting employ two transformers to process spatial and temporal dimensions separately (xu2020spatial; chen2022bidirectional; huo2023hierarchical), while state-of-the-art transformers introduce adaptive local attentions (liu2023spatio), time-delay (jiang2023pdformer), or pre-training (gao2024spatial).
In recent years, simple linear models have been proposed to challenge the effectiveness and efficiency of heavy transformers. DLinear (Zeng2022AreTE) simply separates the input signals as trends and residuals, and employs 2 simple linear layers shared across nodes for prediction, but beats some of the transformer baselines in long-term forecasting. Other representative simple linear models include STID (10.1145/3511808.3557702) and SimpleTM (chen2025simpletm).
@@
A.3 Studies on Directed Graph Laplacians
In the DGSP literature, frequencies of a general graph (directed or undirected) were first defined in Sandryhaila2014 via Jordan decomposition of the adjacency matrix, which is known to be numerically unstable. In Shafipour2019, a set of orthornormal vectors (deemed directed graph frequencies) are computed via a minimization of spectral dispersion in a constrained quadratic programming formulation in eq. (8), which is computationally expensive. In Marques2020, a graph shift operator (GSO) $\mathbf{S}$ is assumed to be diagonalizable into $\mathbf{S}=\mathbf{V}\text{diag}(\{\lambda_{k}\})\mathbf{V}^{-1}$, where the generally non-orthogonal eigenvectors in $\mathbf{V}$ are deemed frequency modes and ordered using a total variation measure in eq. (2), given eigenvalues $\lambda_{k}$’s can be complex-valued. However, asymmetric adjacency and graph Laplacian matrices are not always diagonalizable via the Spectral Theorem. Seifert2021 proposes to augment a directed graph with extra edges to destroy nontrivial Jordan blocks, so that the resulting adjacency matrix becomes diagonalizable. Adding edges, however, fundamentally changes the graph representation of the original pairwise relationships. Chen2023Graph proposes to use singular value decomposition (SVD) on the directed Laplacian to derive frequencies for directed graphs, which is computationally expensive. Kwak2024 proposes to decompose the asymmetric Laplacian matrix into a unitary part (capturing rotation or cycles) and a positive semi-definite part (capturing diffusion-like behavior) also using SVD and polar decomposition, thus suffering the same computation burden. None of these related works proposed variational terms to define directed graph frequencies, which in turn can be simply added to an objective to promote signal smoothness during optimization without first computing frequency components, as we have done.
@@ A.3 Studies on Directed Graph Laplacians
Appendix C Interpreting Linear System Equation 10
@@
${\mathbf{r}}^{\tau}\triangleq\small({\mathbf{L}}^{d}_{r})^{\top}(\frac{{\boldsymbol{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}.$ (55)
Solution ${\mathbf{x}}^{\tau+1}$ to linear system Equation 10 can now be written as
@@
Appendix D Deriving Solution to Optimization Equation 15
The optimization for $\boldsymbol{\phi}$} in objective Equation 15 can be performed entry-by-entry. Specifically, for the $i$-th entry $\phi_{i}$:
@@ Appendix D Deriving Solution to Optimization Equation 15
Alternative solution with subgradients: Denote $g_{1}({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1},g_{2}({\boldsymbol{\phi}})=({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|_{2}^{2}$, the subgradient of $g({\boldsymbol{\phi}})$ is
$\partial g({\boldsymbol{\phi}})=\{{\mathbf{g}}+\nabla g_{2}({\boldsymbol{\phi}})|{\mathbf{g}}\in\partial g_{1}({\boldsymbol{\phi}})\}=\{\mu_{d,1}\sigma+{\boldsymbol{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})|\sigma\in\partial(\|{\boldsymbol{\phi}}\|_{1})\},$ (62)
@@
Appendix E Algorithms for the Unrolled Network
E.1 Parameterizing Conjugated Gradient Method in Solving Linear Equations
@@
Algorithm 1 Conjugate Gradient (CG) Method to Solve ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$ (Unrolled version uCG($\mathbf{A}$}, $\mathbf{b}$}, ${\mathbf{x}}_{0}$))
@@
7:  $\boxed{\alpha_{k}\leftarrow\dfrac{{\mathbf{r}}_{k}^{\top}{\mathbf{r}}_{k}}{{\mathbf{p}}_{k}^{\top}{\mathbf{A}}{\mathbf{p}}_{k}}}$ $\triangleright$ Learnable in unrolling
@@
10:  $\boxed{\beta_{k}\leftarrow\dfrac{{\mathbf{r}}_{k+1}^{\top}{\mathbf{r}}_{k+1}}{{\mathbf{r}}_{k}^{\top}{\mathbf{r}}_{k}}}\triangleright$ Learnable in unrolling
@@
E.2 Detailed Algorithms and Scalability Analysis for the Proposed Model
Algorithms 2 detailedly explained the operations in ADMM blocks corresponding to Figure 3(b) and (c). An overall algorithm of the unrolled network corresponding to Figure 3(a) is shown in Algorithm 3. All matrix multiplication is computed locally to ensure low computational and memory cost with
$({\mathbf{L}}^{u}{\mathbf{x}})_{j}=$ $D_{j,j}^{u}x_{j}-\sum_{i\in{\mathcal{N}}_{j}}W_{i.j}x_{j},\quad({\mathbf{L}}^{u}_{r}{\mathbf{x}})_{j}=x_{j}-\sum_{i\in{\mathcal{N}}_{i}}\frac{W_{i,j}}{\sqrt{D_{i,i}^{u}D_{j,j}^{u}}}x_{i},$ (67)
$({\mathbf{L}}^{d}_{r}{\mathbf{x}})_{j}=$ $x_{j}-\sum_{(i,j)\in{\mathcal{E}}^{d}}\frac{W_{j,i}^{d}}{D_{j,j}^{d}}x_{i},\quad(({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{x}})_{j}=x_{j}-\sum_{(j,i)\in{\mathcal{E}}^{d}}\frac{W_{i,j}^{d}}{D_{i,i}^{d}}x_{i},\quad{\mathcal{L}}^{d}_{r}{\mathbf{x}}=({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}{\mathbf{x}}.$ (68)
Algorithm 2 Forward Propagation of the Unrolled ADMM_Block
1:Initial signal output of the last ADMM block ${\mathbf{x}}\in\mathbb{R}^{T+S+1}$, undirected graph Laplacian matrix ${\mathbf{L}}^{u}$, directed graph Laplacian matrix ${\mathbf{L}}^{d}_{r}$.
2:Network output of the unrolled ADMM block with $M$ layers.
[…]
4:Select real observations with mask: ${\mathbf{y}}\leftarrow{\mathbf{H}}{\mathbf{x}},{\mathbf{y}}\in\mathbb{R}^{T+1}$.
5:for $\tau$ in $0:M$ do $\triangleright$ $M$ ADMM layers
6:  Initialize ${\boldsymbol{\phi}}\leftarrow{\mathbf{L}}^{d}_{r}{\mathbf{x}}$
7:  ${\mathbf{x}}\leftarrow\texttt{uCG}({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho^{\tau}}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}^{\tau}+\rho_{d}^{\tau}}{2}{\mathbf{I}},({\mathbf{L}}^{d}_{r})^{\top}\!(\frac{{\boldsymbol{\gamma}}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}})-\frac{{\boldsymbol{\gamma}}_{u}}{2}+\frac{\rho_{u}^{\tau}}{2}{\mathbf{z}}_{u}-\frac{{\boldsymbol{\gamma}}_{d}}{2}+\frac{\rho_{d}^{\tau}}{2}{\mathbf{z}}_{d}+{\mathbf{H}}^{\top}{\mathbf{y}},{\mathbf{x}})$
8:  ${\mathbf{z}}_{u}\leftarrow\texttt{uCG}(\mu_{u}^{\tau}{\mathbf{L}}^{u}+\frac{\rho_{u}^{\tau}}{2}{\mathbf{I}},\frac{{\boldsymbol{\gamma}}_{u}}{2}+\frac{\rho_{u}^{\tau}}{2}{\mathbf{x}},{\mathbf{z}}_{u})$
9:  ${\mathbf{z}}_{d}\leftarrow\texttt{uCG}(\mu_{d,2}^{\tau}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}^{\tau}}{2}{\mathbf{I}},\frac{{\boldsymbol{\gamma}}_{d}}{2}+\frac{\rho_{d}^{\tau}}{2}{\mathbf{x}},{\mathbf{z}}_{d})\triangleright$ Minimize $\mathbf{x}$} with Equations 10, 11 and 12
10:  ${\boldsymbol{\phi}}\leftarrow\text{soft}_{\mu_{d,1}^{\tau}/\rho^{\tau}}({\mathbf{L}}^{d}_{r}{\mathbf{x}}-{\boldsymbol{\gamma}}/\rho^{\tau})$ $\triangleright$ Minimize $\phi$ with Equation 16
11:  ${\boldsymbol{\gamma}}\leftarrow{\boldsymbol{\gamma}}+\rho^{\tau}({\boldsymbol{\phi}}^{\tau+1}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})$
12:  ${\boldsymbol{\gamma}}_{u}\leftarrow{\boldsymbol{\gamma}}_{u}+\rho_{u}^{\tau}({\mathbf{x}}-{\mathbf{z}}_{u})$
13:  ${\boldsymbol{\gamma}}_{d}\leftarrow{\boldsymbol{\gamma}}_{d}+\rho_{d}^{\tau}({\mathbf{x}}-{\mathbf{z}}_{d})$ $\triangleright$ Update multipliers with Equations 17 and 18
14:end for
15:return $\mathbf{x}$}
Algorithm 3 Overall Algorithm of the Proposed Unrolled Network
1:An observation from the training set ${\mathbf{y}}\in\mathbb{R}^{T+1}$.
2:The reconstructed signal ${\mathbf{x}}\in\mathbb{R}^{T+S+1}$ containing both the obseration and prediction.
[…]
7:for $i$ in $0:B$ do $\triangleright$ $B$ ADMM blocks and graph learning modules
8:  ${\mathbf{e}}_{i}\leftarrow[{\mathbf{x}}_{i};{\mathbf{e}}_{\text{st},i}],{\mathbf{f}}_{i}\leftarrow F^{u}({\mathbf{e}}_{i})$ for $i=1,\dots,N$ $\triangleright$ Feature extraction (Appendix F.3.2)
9:  ${\mathbf{L}}^{u(h)}\leftarrow\texttt{UGL}_{h}({\mathbf{f}}_{1},\dots,{\mathbf{f}}_{N}),\quad{\mathbf{L}}^{d(h)}_{r}\leftarrow\texttt{DGL}_{h}({\mathbf{f}}_{1},\dots,{\mathbf{f}}_{N}),h=1,2,\dots,H$
[…]
11:  ${\mathbf{x}}_{\text{new}}^{(h)}\leftarrow\texttt{ADMM\_Block}_{h}({\mathbf{x}},{\mathbf{L}}^{u(h)},{\mathbf{L}}^{d(h)}_{r}),h=1,2,\dots,H$
12:  ${\mathbf{x}}_{\text{new}}=\sum_{h=1}^{H}a_{h}{\mathbf{x}}_{\text{new}}^{(h)}$ $\triangleright$ Learnable merge of multi-head results
13:  ${\mathbf{x}}\leftarrow p_{i}{\mathbf{x}}_{\text{new}}+(1-p_{i}){\mathbf{x}}$ $\triangleright$ Residual connections with learnable parameter $p_{i}$
14:end for
15:return $\mathbf{x}$}
E.3 Scability Analysis
our model scales linearly as long as the city map is sparse. The conjugated gradient algorithm is detailed in Appendix E, which involves only sparse matrix-vector multiplications in each iteration. Specifically, each matrix multiplication involves graph Laplacians describing pairwise spatial and temporal similarities in local neighborhoods. Thus, as long as the temporal window $W$ and number of neighbors $k$ do not change, the computation cost in unrolled CG grows only linearly with the increase of total time steps and number of nodes. Our graph learning modules also calculate and save the edge weights locally in each node’s neighborhood, as is demonstrated in Equation 68. Taken together, with fixed $W$ and $k$, the size of our model, computation time, and memory cost all grow linearly with the total time steps and network size.